OpenStudy (anonymous):

Use the principle of induction to prove that the number of derangements of n,dn, can be given by the recursion formula d1 = 0, d2 = 1, dn = (n-1)(dn-1 + dn-2) (n >= 3).

jhonyy9 (jhonyy9):

first make n=k is true and after you need to prove that this is true for n=k+1 too

jhonyy9 (jhonyy9):

d3=(3-1)(d2 +d1) =2*(1+0) =2

OpenStudy (anonymous):

1st take like this n=0,1

OpenStudy (anonymous):

i mean d1=(0-1)(d1-1+d2-2)

OpenStudy (anonymous):

here we can solve like this (-1)(d1-1+d1-2) =-d1+1+d1-2 =-1 and also d2=1,so,d2=(1-1)(d2-1+d2-1) =0 so the answer is (0,-1)

OpenStudy (anonymous):

oh sorry mistake

OpenStudy (anonymous):

u ca take opposite as n=k so here we can solve also like this i mean do the opposite then like n=K+1 =0-1 =-1 n=k+1 =-1+1 =0

OpenStudy (anonymous):

whats ur answer?

jhonyy9 (jhonyy9):

,,down12" check it please newly because when n=1 so than d1=(1-1)(---) =0*(---) =0 -so because we know that one indifferent number time zero will be equal zero yes ?